OpenStudy (anonymous):

Log question, 10^(x+3) = 6^(2x), solve for x

OpenStudy (anonymous):

the answer is \[\frac{ 3\ln10 }{ \ln18-\ln5 }\] but i don't know how to get there, help please?

OpenStudy (phi):

use the rule \[ ln(x^a) = a \cdot ln(x) \]

OpenStudy (anonymous):

yea i tried, got \[\frac{ 2x \ln6 }{ lnl0 } - x = 3\] but dunno how to continue..

OpenStudy (phi):

(x+3) ln(10)= 2x ln(6) is how I start out

OpenStudy (anonymous):

yea but Im unable to get the answer exactly like that

OpenStudy (phi):

continuing (x+3) ln(10)= 2x ln(6) rewrite 2 x ln(6) as x ln(6^2) because the final answer seems to want this also distribute the ln(10) on the left side: x ln(10) + 3 ln(10)= x ln(36) they have 3 ln(10) up top, so subtract x ln(10) from both sides

OpenStudy (anonymous):

what about x= 3/2log(6)-1

OpenStudy (anonymous):

would that be right?

OpenStudy (phi):

x ln(36)- x ln(10) = 3 ln(10) x (ln(36)- ln(10)) = 3 ln(10) to simplify the left side. notice that ln(36)- ln(10) = ln(36/10)= ln(18/5)= ln(18)-ln(5) x ( ln(18)-ln(5))= 3 ln(10)

OpenStudy (phi):

divide through to get their answer x= 3 ln(10)/ ( ln(18)- ln(5)) any of the steps confusing?

OpenStudy (phi):

@KE8717504 Yes, if you mean log base 10, and you put parens in like this: 3/(2 log(6)-1)

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

OMFG @phi THANKS SO MUCH

OpenStudy (phi):

If we pick up where you left off: yea i tried, got \[\frac{ 2x \ln6 }{ ln10 } - x = 3 \] but dunno how to continue.. factor out the x on the left side. also bring the back inside the ln: \[ (\frac{ln36}{ln(10) }-1)x= 3 \] or \[ (\frac{ln36}{ln10 }-\frac{ln10}{ln{10}})x= 3 \] multiply both sides by ln10 \[(ln36-ln10)x= 3 ln10 \] simplify the left side like before, and you get to the answer...